#2 g and #3&4 please!!!#2 For each of the following, first sketch (roughly) the region enclosed by the given curves, labelling any important points
Answer #1

g> y= am x+y=3, xxo y=aur 4 yazné m ya (6,39 Al van X81212 (3,09 Telo,u kry=3 We have to find A and B. ary: -3 » M= (3-4) y =area. Now we have to find the 2. L J As - -)a 2 G 씨 AC). 13 1 2 ( [국] - -- + { 황 사 --2- . - 3 + 122. C 3 아 17 7 Aus)